FLUORESCENCE SPECTROSCOPY OF URANIUM FLUORIDE COMPOUNDS IN SOLID Ar

\underline{JIN JIN}, CHRISTOPHER J. LUE, JONATHAN C. RIENSTRA-KIRACOFE, AND MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

Spectroscopic studies of uranium fluorides are being carried out to provide benchmark data for evaluating relativistic quantum chemical calculations. Matrix isolation techniques are being used to obtain survey spectra for a variety of UF_{n} (n≤ 6) species. Initial experiments have focused on UF_{6} and fragments produced by passing UF_{6}/Ar mixtures through a microwave discharge.

UF_{6}/Ar matrices were examined using 355 and 266 nm excitation. The former excites the A-X band, producing a longlived (600μs) emission that spans the 408-435 nm range. This system has been characterized previously^{a}. Excitation at 266 nm produced a new emission system in the 313-417 nm range with an upper state lifetime of 500 ns. This spectrum was dominated by a nearly harmonic progression with a frequency of 670 cm^{−1}. As 266 nm light is known to dissociate gas phase UF_{6}, the nature of the emitting state is in question.

Passage of the UF_{6}/Ar mixture through a microwave discharge prior to matrix deposition destroyed all of the hexafluoride (the matrices did not show the A-X or new emission bands described above). With 355 nm excitation a new system of emission bands was observed in the 427-500 nm range. These emissions are consistent with emissions from UF and/or UF_{2}.

Experiments to identify the carriers of the new band systems are in progress. A summary of the results will be presented.

aJ. C. Miller, S. W. Allison and L. Andrews, J. Chem. Phys. 70, 3524 (1979).